OpenStudy (anonymous):

Solve the triangle. B = 36°, a = 38, c = 18 So far I have gotten that b=25.7. But now I'm stuck.

OpenStudy (anonymous):

then do sin theorem a/sin(A)=b/sin(B)=c/sin(C)

OpenStudy (anonymous):

I got 54.97858039 as A. That is not one of the possible answers so I know I did something wrong. I took b/sinB, 38/sinANS

OpenStudy (phi):

I just used your numbers, but got something closer to 60 deg

OpenStudy (anonymous):

Hmm. Something is wrong then. These are the possible answers: b ≈ 41.7, C ≈ 28.2, A ≈ 115.8 b ≈ 25.7, C ≈ 24.2, A ≈ 119.8 b ≈ 25.7, C ≈ 116.8, A ≈ 27.2 b ≈ 41.7, C ≈ 24.2, A ≈ 119.8

OpenStudy (phi):

it looks like you used 36 instead of 38 for the length of a

OpenStudy (phi):

it is clearly choice B or C A= 60.3, or If the triangle is obtuse, A= 180-60.3= 119.7

OpenStudy (anonymous):

So if I get an answer and its not there, try subtracting it from 180? How do I know if its obtuse?

OpenStudy (phi):

Law of sines can be ambiguous.

OpenStudy (phi):

one clue is that the biggest side goes with the biggest angle. if one side is "huge" compared to the others, look out for obtuse.